RHIC pp Set-Up: Beams are now in the center of the beam pipe, after RF adjustments of ~40Hz corrected for radii that were off by ~1mm.
Number of bunches increased to 37 with Ramp efficiency up to 90%. There were some first collisions overnight. Blue tunes are at 0.93/0.94 (instead of 0.94/0.93) Blue beam loss shoot up every minute or so.
There was a snake quench this morning. It occurred when the loss monitors showed a hit at the snake, even through the AtR switching magnet was off. There was a discussion of how this could happen, or even whether it could happen.
STAR does not have its ZDCs turned on, so we can't see collisions at STAR.
For today, machine development is to swap the Blue tunes. Todd should look into the Yellow optics. Vadim will continue working on orbits at the IPs, and separation bumps. Waldo will start rotator ramp development.
This evening, we should measure Blue profile with Polarimeter; verify injection matching improvement. Verify the fill pattern at store. Measure polarization at store. Haixin reports that currently the polarization at extraction from the AGS is 60% at 1E11, dropping to 50% at 1.5E11.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |